OpenStudy (anonymous):

41% of 78 is what? 41% of 78 is what? @Mathematics

OpenStudy (turingtest):

0.41x78=?????????

OpenStudy (anonymous):

thank you! & do you now how to make a proportion to that? :)

OpenStudy (anonymous):

okay! thank you<3

OpenStudy (anonymous):

and like do you know when to make what proportion? like for this problem: What is 12% of 17.5?

OpenStudy (turingtest):

I'm not totally sure what you mean by "make a proportion" but to get some kind of fraction just remember A% is A/100 or A% of B is Bx(A/100) so 12% of 17.5 is 17.5(12/100)=210/100=21/10

OpenStudy (turingtest):

stupid server... :/

OpenStudy (anonymous):

i'll like draw something now. gonna take like 3 mins..

OpenStudy (anonymous):

 this is like; what percent of 29 is 3?